Suppose 1900 J of heat energy are put into a sample of liquid water at 21°C, and the final temperature of the water is 76°C. How many grams of water are present?
1
Expert's answer
2012-10-10T08:41:11-0400
Heat capacity for water is C= 4218 J/Kg*K (it's for delta T =1 and 1 kg ) we
have delta T=55 and x kg
For 1 kg and delta T = 55 you need 55*4218 J=231990
j
For x kg and delta T = 55 we have 1900 J

x=1*1900/231990=0.0082 kg
or 8.2 ml